OpenStudy (anonymous):

x^2+4y^2-16=0 write the standard form of this equation

OpenStudy (campbell_st):

its an ellipse so its \[\frac{ax^2}{ab} + \frac{by^2}{ab} = r^2\] so a = 1 and b = 4 then simplify after that

OpenStudy (anonymous):

thank you